OpenStudy (anonymous):

A kite rises at an angle of 11° with the ground. Find the distance the kite has flown (hint: the length of the string) when it has covered a horizontal distance of 100 feet. Round your answer to the nearest foot.

OpenStudy (anonymous):

98 feet 102 feet 514 feet 524 feet

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

 Using trig we have \[\cos(11)=\frac{ 100 }{ x }\] so \[x = \frac{ 100 }{ \cos(11) }\] \[x \approx 102\]